OpenStudy (anonymous):

Find the solution of y=6x+6 for x=-2 WILL GIVE MEDAL FOR CORRECT ANSWER!

OpenStudy (mathstudent55):

Replace x with -2 and calculate.

OpenStudy (johnweldon1993):

The "x = -2" just tells you to make every 'x' you see In your equation a -2 In your case y = 6x + 6 will become y = 6(-2) + 6 Now calculate

OpenStudy (anonymous):

-2,-6

OpenStudy (johnweldon1993):

So you have the y = 6(-2) + 6 so what is 6 times -2? -12 so now it's y = -12 + 6 y = ....? *Since you gave a medal I'm assuming you already have the answer*